INAF LBC+MODS queue observing, 2011 October 22-23

Observer: K. Boutsia, M. Fumana, P. Santini
Telescope Operator: G. Bechetti
Support Astronomer: D. Thompson
Instrument: MODS + LBC

Summary

Clear sky in the afternoon and throughout the night. Most probably photometric .
We started with MODS and switched to LBC. No problems with MODS tonight.
LBC has given some problemst close to the dawn. The 1st chip of ther red channel has shown stripes. This defect seemed very similar to what happened with the blue channel on 20111020.
The error LBC was giving on the red side in the LogAnalyzer was camera system encountered a problem .
We turn OFF/ON the LBC, but then we could not upload OBs anymore. We did lbckill - lbcstart, and still we could not upload the OBs.
After some digging we found out that the OB upload error was due to the LBC disk being full. We transfered some log files to make space and after that LBC has worked fine.
Since we did both: turn OFF/ON and lbckill-lbcstart, we don't know exactly which operation has solved the 'stripes' problem on the red side.
35min were lost because of the LBC error.

Data Summary

Project Objects Filter ExpTime Notes

Nonino a2261 Clear - GG495 20min seeing ~0.9, sources probably not resolved
Clementini AndXXV Bbess - Vbess 2hr blue image not well collimated in one OB
Clementini AndXXVII Bbess - Vbess 2hr blue image not well collimated in one OB
STD Stetson_L92 Bbess - Vbess / Ibess   1 position - airmass 1.54
Lardo NGC2419 Uspec - Vbess / Ibess 1hr  
Bragaglia NGC2225 Bbess - Vbess / Ibess 20min done at the beginning of twilight
STD Stetson_L92 Uspec / Bbess - Vbess / Ibess   4 positions - airmass ~1.2
STD SA98 Bbess - Vbess / Ibess   4 positions - airmass ~ 1.2

Details

23:48 MODS calibrations - some of them need to be repeated

01:10 Pointing correction

01:30 Still too bright for the WFSC

01:38 Preset to a2261_pa0.acq
  • seeing on the guider: 1.1 - 1.4
  • seeing on th guider after picking up the guiding star of Nonino: 1.0 - 1.1
  • mods1r.20111023.0024.fits - 600sec

01:48 execMODS a2261_pa0.obs

02:06 Preset to a2261_pa90.acq
  • seeing on the guider: 1.4

02:14 execMODS a2261_pa90.obs
  • repeat a2261_pa90.obs because we can see the spectrum and we have some time
  • guider seeing 1.2
  • abort the second exposure because we can not resolve the three sources
  • mods1r.20111023.0028.fits - 600sec

02:31 Preset to feige110.acq (standard found in the MODS calibration directory)

02:42 execMODS feige110.obs (1'' slit and 5'' slit)
  • mods1r.20111023.0031.fits - 120sec
  • mods1r.20111023.0032.fits - 30sec

02:52 SWITCH TO LBC

03:13 Slew to pointing star WT10_206
  • dofpia,/first (03:16) (only 3 or 4 iterations to converge)

03:18 lbcrangebal on WT10_206

03:26 Slew to Stetson_L92_5 for standard calibration (airmass=1.54)

03:32 Slew to M31_AndXXV_400s_18
  • dofpia

03:40 M31_AndXXV_400s_18.ob
  • seeing measured on first images ~1.24(blue) ~0.87 (red)
  • seeing measured on images ~1.24(blue) ~0.79 (red)
  • blue profiles are slightly elongated
  • seeing measured on last images ~1.24 (blue) <0.88 (red)

04:02 Slew to M31_AndXXVII_400s_18
  • dofpia

04:06 M31_AndXXVII_400s_18.ob
  • seeing measured on first images ~1.13 (blue) ~0.68 (red)
  • seeing measured on images ~1.13 (blue) ~ 0.72 (red)
  • blue image is not well collimated

04:29 Slew to M31_AndXXV_400s_19
  • dofpia
  • dofpia (04:33)

04:35 M31_AndXXV_400s_19.ob
  • seeing measured on first images ~0.95 (blue) ~0.79 (red)
  • seeing measured on images ~1.0 (blue) ~ 0.8 (red)
  • seeing measured on images ~0.9 (blue) ~ 0.8 (red)

04:57 Slew to M31_AndXXVII_400s_19
  • dofpia

05:02 M31_AndXXVII_400s_19.ob
  • seeing measured on first images ~ 0.9 (blue) ~ 0.7 (red)
  • seeing measured on images ~ 1.0 (blue) ~ 0.9 (red)

05:28 Slew to M31_AndXXV_400s_20
  • dofpia

05:31 M31_AndXXV_400s_20.ob
  • seeing measured on images ~1.0 (blue), <1.0 (red)

05:55 Slew to M31_AndXXVII_400s_20
  • dofpia

05:59 M31_AndXXVII_400s_20.ob
  • seeing measured on images ~0.8 (blue), 0.7 (red)

06:21 Stetson_L92_1 (airmass ~1.2)
  • it has been altered to include Uspec filter - 4pos

06:42 Slew to M31_AndXXV_400s_21
  • dofpia

06:54 M31_AndXXV_400s_21.ob
  • seeing measured on images ~0.9 (blue), ~0.95 (red)

07:16 Slew to M31_AndXXVII_400s_21
  • dofpia

07:22 M31_AndXXVII_400s_21.ob
  • seeing measured on images <0.88 (blue), < 0.97 (red)
  • seeing measured on images <0.9

07:45 Slew to M31_AndXXV_400s_22
  • dofpia

07:50 M31_AndXXV_400s_22.ob
  • seeing measured on images <0.7''

08:12 Slew to M31_AndXXVII_400s_22
  • dofpia

08:16 M31_AndXXVII_400s_22.ob
  • seeing measured on images

08:39 Slew to M31_AndXXV_400s_23
  • dofpia

08:43 M31_AndXXV_400s_23.ob
  • seeing measured on images ~0.6''

09:06 Slew to M31_AndXXVII_400s_23
  • dofpia

09:10 M31_AndXXVII_400s_23.ob
  • seeing measured on images <0.79
  • seeing measured on images ~0.75 (blue) <0.9 (red)
  • seeing measured on images ~0.7 (blue) ~0.67 (red)

09:34 Slew to N2419_1
  • dofpia

09:49 N2419_1.ob
  • seeing measured on images <0.8 (blue) ~0.86 (red)
  • seeing measured on images <0.8 (blue) ~0.56 (red)

09:58 Slew to N2419_2
  • dofpia

10:05 N2419_2.ob
  • seeing measured on images <0.72 (red)
  • seeing measured on images <0.89 (blue and red)
  • dofpia

10:27 N2419_3.ob
  • seeing measured on images ~0.9
  • dofpia

10:50 N2419_4.ob
  • seeing measured on images ~0.9 (blue) ~0.8 (red)

10:57 N2419_5.ob
  • seeing measured on images ~0.86 (blue) ~0.8 (red)

11:03 N2419_6.ob
  • seeing measured on images ~0.68 (blue) ~0.6 (red)

11:13 Slew to ngc2225
  • dofpia

11:22 ngc2225.ob
  • OB interrupted after the first image - because chip1 of red channel has saturated bands it has the same defect

11:25 repeat OB - the chip has still stripes...
  • interrupt OB - take a fast image to see if it persists
  • from the LOG Analyzer:
camera system encountered a problem 
OB execution error: none [src/lbc/actions.c:1155] 
  • When we tried to turn OFF and turn ON - we got the error - LBC restarted after a crash
  • We cannot upload the OBs
Warning: move_uploaded_file() [function.move-uploaded-file]: Unable to move '/tmp/phpP5F8Kc' to '/home/lbccontrol/uploads/import.ob' in /home/lbccontrol_binocular_orig/www/ui/3.php on line 45
  • Dave runs lbckill.sh - lbcstart.sh
  • the problem with upoloading the OBs was caused because LBC/home was FULL - Dave had to move the *log file to some other directories to make space.

12:00 try the preset again - ngc2225
  • dofpia
  • in the first image we do not see the defect anymore... it is probably the lbckill-lbcstart that solved the camera problem
  • seeing on the images: ~0.9 (blue) - 0.8 (red)
  • airmass = 1.365
  • seeing on the images: <0.7-0.8
  • seeing on the images: ~0.79 (blue) - 0.7 (red)
  • airmass = 1.371
  • done at the beginning of twilight

12:19 Test SkyFlats
  • Uspecial - Vbessel (1sec-1sec) ~500 counts in both channels
  • wait a few minutes, sky is still too dark
  • Uspecial - Vbessel (1sec-1sec) - scale 10 (12:25) ~900 counts in both channels
  • too early for flats

12:27 Standard calibration: SA98 (airmass~1.2)
  • dithering 4 positions

12:34 Test SkyFlats
  • Uspecial - Vbessel (1sec-1sec) ~600 counts in both channels
  • Uspecial - Vbessel (1sec-1sec) ~800 counts in blue, ~700 counts in red (12:41)
  • stars are still visible
  • Uspecial - Vbessel (1sec-1sec) ~1100 counts in blue, ~900 counts in red (12:44)

12:47 SkyFlats
  • Uspecial - Vbessel (1sec-1sec) - scale 10 ~12000 counts in blue, ~7000 counts in red (12:44)
  • Uspecial - Vbessel (1sec-1sec) - scale 10 ~16000 counts in blue, ~12000 counts in red (12:51)
  • Uspecial - Vbessel (1sec-1sec) - scale 10 ~27000 counts in blue, ~16000 counts in red (12:54)
  • 180 degree
  • Uspecial - Vbessel (1sec-1sec) - scale 7 ~33000 counts in blue, ~20000 counts in red (12:56)
  • Uspecial - Vbessel (1sec-1sec) - scale 7 ~42000 counts in blue, ~26000 counts in red (12:58)
  • Uspecial - Vbessel (1sec-1sec) - scale 2 ~38000 counts in blue, ~30000 counts in red (13:00)

01:07 start skyflat Bbessel - Ibessel but it is too late! - END OF NIGHT
  • continue MODS calibrations for the LS at the beginning of the night.
  • Imaging flats (030-034) Blue, (038-042) red
  • LS60x5 flat (036-040) Blue, (44-48) red
  • LS60x5x1.0 flat (042-046) Blue, (50-54) red

-- KonstantinaBoutsia - 23 Oct 2011
Topic revision: r4 - 26 Nov 2011, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback